# BEST-STD 2.0: Balanced and Efficient Speech Tokenizer for Spoken Term Detection

### v1

@@ 1 Introduction
The rapid growth of spoken content across digital platforms has underscored the need for robust and efficient Query-by-Example Spoken Term Detection (QbE-STD) systems. However, existing systems often struggle in noisy and reverberant environments, limiting their effectiveness in real-world applications such as voice search [wang2008introduction].
@@
The ASR-based approaches [mamou2007vocabulary, miller2007rapid, wang2008comparison] represent speech using subword lattices to handle OOV terms but require highly accurate ASR models, which are difficult to train for short, low-context queries. Direct audio comparison systems [ram2018cnn, ram2020neural] circumvent transcription by performing direct audio-template matching using Segmental Dynamic Time Warping (SDTW) [tsai2021segmental], but the high computational cost of SDTW hinders scalability. Other approaches [he2016multi, chung2016audio, kamper2016deep] learn discriminative acoustic word embeddings, but these methods require precise word boundaries during training and inference, making them impractical for most real-world scenarios. Moreover, most existing STD systems are designed for clean acoustic environments and suffer severe degradation under noise and reverberation.
@@
To alleviate these issues, BEST-STD [singh2025best] introduced a discrete tokenization strategy that converts speech into subword-like units, enabling fast retrieval with text-based search algorithms, eliminating boundary annotations at inference, and efficiently handling OOV terms. However, this approach still produces tokens that are sensitive to acoustic conditions and exhibit low entropy, limiting both robustness and discriminability. Existing speech tokenizers [hsu2021hubert, chen2022wavlm, zhang2023speechtokenizer] that generate semantic tokens also remain entangled with speaker characteristics and other acoustic cues.
@@
This paper introduces BEST-STD 2.0, a novel speech tokenizer designed to produce speaker-agnostic and noise-robust tokens. Our framework transforms input speech into a contextual frame-level embedding sequence using a bidirectional Mamba encoder [singh2025best], followed by tokenization through vector quantization [van2017neural]. To ensure consistent tokenization, we employ a self-supervised training objective that maps different utterances of the same term to the same token sequence, even under distortions like noise and reverberation. Furthermore, to address the common issue of token index collapse [huh2023straightening] in discrete representation learning, we reformulate token learning as a balanced clustering problem and introduce an optimal-transport regularization [cuturi2013sinkhorn] that promotes uniform codebook usage and enhances discriminability. Empirical evaluations on the LibriSpeech and TIMIT datasets with added real-world noise scenarios show that our system outperforms existing STD baselines across a range of acoustic conditions, demonstrating improved robustness and search efficiency. By enabling text-like search capabilities over raw speech, our method provides a scalable solution for spoken content retrieval. Overall, the main contributions of our work are as follows:
@@
  • A TF-IDF–based retrieval strategy enabling fast and efficient spoken-term retrieval.
@@ 2.1 Model Architecture
We adopt the design choices for the Mamba feature encoder as described in [singh2025best]. The encoder $f:X\rightarrow Z$ comprises $L$ layers of bidirectional Mamba blocks, which map the input audio feature sequence $X=\{x_{1},\dots,x_{t}\}$ to the corresponding contextual speech representations $Z=\{z_{1},\dots,z_{t}\}$. The output sequence embeddings of the final encoder layer are projected into a $d$-dimensional space, followed by $L_{2}$ normalization to ensure a unit norm.
@@ 2.2 Self-Supervised Learning Framework
The above formulation can be interpreted as a case of the optimal transport (OT) problem, which can be effectively computed in almost linear time with the Sinkhorn-Knopp algorithm [cuturi2013sinkhorn]. In the OT framework, we define the cost of assigning the $t^{th}$ sample $z_{t}$ to $c_{k}$ as -$s_{k}(z_{t})$. The resulting solution provides $p(z_{t}|c_{k})$, which we use in Eq. 7.
@@ 2.3 Indexing and Retrieval
Given a set $\mathcal{D}$ of audio tracks, each track $a_{i}\in\mathcal{D}$ is divided into overlapping segments of length $l$ s with a hop size of $h$ s. For each segment, we compute its representation $Z=\{z_{t}\}_{t=1}^{T}$, followed by its corresponding tokenized representation $Z^{\prime}=\{i_{t}\}_{t=1}^{T}$, where each $i_{t}$ denotes the index of the nearest codeword for $z_{t}$ in the codebook $C$, with $i_{t}\in\{1,2,\dots,K\}$. Subsequently, we construct a TF-IDF representation for each tokenized sequence $Z^{\prime}$ in the database and index them using IVF-PQ [jegou2010product] for fast retrieval.
@@ 3.1 Databases
We trained the model on the LibriSpeech train-clean-360 subset and used the test-clean subset for validation. To construct the speech archive and queries, we selected the train-clean-100 subset, which contains approximately 100 hours of spoken content. This setup ensures that the evaluation is always conducted on speakers unseen during training our model. To assess cross-dataset generalization, we also evaluated our model on the TIMIT dataset [garofolo1993darpa]. We employed Montreal Forced Aligner [mcauliffe2017montreal] to obtain word alignments for both datasets. To emulate realistic acoustic conditions, the training data were augmented with a diverse range of background noise types and reverberation effects by incorporating noise samples and room impulse responses (RIRs) from the MUSAN corpus [snyder2015musan]. For evaluation, we employed RIRs from the Aachen Impulse Response Database [jeub2009binaural] and noise clips from the ETSI background noise database [transmission1994speech], both of which contain recordings of real-world acoustic environments such as offices, public spaces, and street settings.
@@ 3.2 Evaluation
We evaluated the spoken term detection baselines using the Mean Term Weighted Value (MTWV) [shah2020cross] as the primary performance metric. Also, to assess the robustness of speech tokenizers against speaker variations and audio distortions, we measured the Jaccard similarity [liu2009learning] between the tokenized representations of different utterances of the same spoken term.
@@ 3.3 Baselines
We evaluated ASR-based baselines for STD tasks, including HuBERT [hsu2021hubert] and WavLM [chen2022wavlm], by extracting posterior tokens from these models. Also, we evaluated speech tokenizers that generate semantic speech tokens, such as WavLM, SpeechTokenizer [zhang2023speechtokenizer], and BEST-STD [singh2025best], for STD tasks. For WavLM, we used the pretrained KMeans model available in the SpeechBrain toolkit [ravanelli2021speechbrain], which was trained on features extracted from the 23rd layer of the large variant of WavLM. For SpeechTokenizer [zhang2023speechtokenizer], we extracted the semantic tokens generated by the first quantizer in its residual vector quantization (RVQ) stack.
@@ 3.4 Implementation details
HuBERT-Large [hsu2021hubert] 32 0.73 0.46 0.59 0.67 0.71 0.72 0.24 0.37 0.49 0.58 0.64
WavLM-Large [chen2022wavlm] 32 0.72 0.62 0.67 0.70 0.71 0.71 0.52 0.60 0.65 0.68 0.70
@@
SpeechTokenizer [zhang2023speechtokenizer] 1024 0.45 0.09 0.12 0.15 0.18 0.19 0.03 0.04 0.05 0.07 0.08
WavLM-Large [chen2022wavlm] 1000 0.40 0.18 0.19 0.21 0.22 0.23 0.16 0.17 0.18 0.18 0.20
BEST-STD [singh2025best] 1024 0.72 0.21 0.29 0.42 0.60 0.65 0.19 0.22 0.38 0.55 0.62
@@ IV OOV IV OOV
HuBERT-Large [hsu2021hubert] 0.13 0.21 0.30 0.40 0.47 0.47 0.16 0.27 0.34 0.40 0.41 0.43 0.14 0.22 0.31 0.43 0.49 0.51 0.16 0.28 0.37 0.43 0.44 0.46
WavLM-Large [chen2022wavlm] 0.31 0.36 0.43 0.52 0.55 0.58 0.29 0.37 0.41 0.42 0.43 0.45 0.33 0.35 0.44 0.52 0.55 0.61 0.33 0.41 0.46 0.47 0.49 0.50
@@
SpeechTokenizer [zhang2023speechtokenizer] 0.14 0.27 0.39 0.49 0.52 0.53 0.13 0.21 0.30 0.42 0.48 0.49 0.15 0.28 0.42 0.53 0.56 0.57 0.15 0.26 0.34 0.43 0.48 0.52
WavLM-Large [chen2022wavlm] 0.17 0.34 0.40 0.53 0.55 0.55 0.17 0.25 0.35 0.43 0.47 0.49 0.19 0.38 0.44 0.57 0.59 0.61 0.19 0.29 0.35 0.46 0.47 0.51
BEST-STD [singh2025best] 0.27 0.35 0.43 0.50 0.57 0.62 0.22 0.29 0.37 0.44 0.49 0.54 0.29 0.38 0.47 0.54 0.62 0.66 0.25 0.33 0.40 0.49 0.50 0.56
@@
HuBERT-Large [hsu2021hubert] 0.02 0.06 0.09 0.13 0.21 0.24 0.02 0.07 0.12 0.20 0.26 0.29 0.03 0.08 0.12 0.23 0.25 0.27 0.08 0.15 0.24 0.26 0.28 0.30
WavLM-Large [chen2022wavlm] 0.11 0.18 0.24 0.30 0.32 0.36 0.15 0.22 0.29 0.31 0.35 0.37 0.12 0.21 0.23 0.35 0.37 0.39 0.18 0.24 0.31 0.35 0.39 0.41
@@
SpeechTokenizer [zhang2023speechtokenizer] 0.03 0.05 0.11 0.14 0.18 0.20 0.02 0.04 0.06 0.11 0.13 0.16 0.05 0.12 0.18 0.19 0.23 0.23 0.07 0.11 0.14 0.18 0.21 0.23
WavLM-Large [chen2022wavlm] 0.06 0.12 0.19 0.25 0.34 0.39 0.04 0.07 0.14 0.21 0.27 0.31 0.08 0.16 0.23 0.26 0.30 0.36 0.10 0.17 0.23 0.25 0.29 0.30
BEST-STD [singh2025best] 0.18 0.26 0.34 0.40 0.46 0.51 0.13 0.20 0.27 0.33 0.39 0.43 0.20 0.28 0.36 0.44 0.49 0.54 0.17 0.26 0.33 0.34 0.42 0.48
@@ 4.2 Codebook analysis
As shown in Table 3, our approach consistently achieves a normalized entropy close to 1, indicating near-perfect balance across different codebook sizes, ranging from 1024 to 4096. In contrast, KMeans-based tokenization [singh2025best] yields substantially lower entropy that drops further as the codebook size increases. Furthermore, we encountered codebook collapse when employing a trainable codebook regularized with KL divergence [baevski2020wav2vec], despite extensive optimization efforts. These results demonstrate that our approach effectively overcomes the codebook collapse problem, enabling stable and scalable learning of large speech-token codebooks.
@@
KL Divergence [baevski2020wav2vec] Learnable 0.63 0.50 0.38
BEST-STD [singh2025best] Kmeans 0.76 0.65 0.43

### v2

@@ 1 Introduction
The rapid growth of spoken content across digital platforms has underscored the need for robust and efficient Query-by-Example Spoken Term Detection (QbE-STD) systems. However, existing systems often struggle in noisy and reverberant environments, limiting their effectiveness in real-world applications such as voice search.
@@
The ASR-based approaches [13, 15, 26] represent speech using subword lattices to handle OOV terms but require highly accurate ASR models, which are difficult to train for short, low-context queries. Direct audio comparison systems [16, 17] circumvent transcription by performing direct audio-template matching using Segmental Dynamic Time Warping (SDTW) [24], but the high computational cost of SDTW hinders scalability. Other approaches [6, 3, 11] learn discriminative acoustic word embeddings, but these methods require precise word boundaries during training and inference, making them impractical for most real-world scenarios. Moreover, most existing STD systems are designed for clean acoustic environments and suffer severe degradation under noise and reverberation.
@@
To alleviate these issues, our previous work, BEST-STD [20], introduced a discrete tokenization strategy that converts speech into subword-like units, enabling fast retrieval with text-based search algorithms, eliminating boundary annotations at inference, and efficiently handling OOV terms. However, this approach still produces tokens that are sensitive to acoustic conditions and exhibit low entropy, limiting both robustness and discriminability. Existing speech tokenizers [7, 2, 27] that generate semantic tokens also remain entangled with speaker characteristics and other acoustic cues.
@@
This paper introduces BEST-STD 2.0, a novel speech tokenizer designed to produce speaker-agnostic and noise-robust tokens. Our framework transforms input speech into a contextual frame-level embedding sequence using a bidirectional Mamba encoder [20], followed by tokenization through vector quantization [25]. To ensure consistent tokenization, we employ a self-supervised training objective that maps different utterances of the same term to the same token sequence, even under distortions like noise and reverberation. Furthermore, to address the common issue of token index collapse [8] in discrete representation learning, we reformulate token learning as a balanced clustering problem and introduce an optimal-transport regularization [4] that promotes uniform codebook usage and enhances discriminability. Empirical evaluations on the LibriSpeech and TIMIT datasets with added real-world noise scenarios show that our system outperforms existing STD baselines across a range of acoustic conditions, demonstrating improved robustness and search efficiency. By enabling text-like search capabilities over raw speech, our method provides a scalable solution for spoken content retrieval. Overall, the main contributions of our work are as follows:
@@ 2.1 Model Architecture
We adopt the design choices for the Mamba feature encoder as described in [20]. The encoder $f:X\rightarrow Z$ comprises $L$ layers of bidirectional Mamba blocks, which map the input audio feature sequence $X=\{x_{1},\dots,x_{t}\}$ to the corresponding contextual speech representations $Z=\{z_{1},\dots,z_{t}\}$. The output sequence embeddings of the final encoder layer are projected into a $d$-dimensional space, followed by $L_{2}$ normalization to ensure a unit norm.
@@ 2.2 Self-Supervised Learning Framework
The above formulation can be interpreted as a case of the optimal transport (OT) problem, which can be effectively computed in almost linear time with the Sinkhorn-Knopp algorithm [4]. In the OT framework, we define the cost of assigning the $t^{th}$ sample $z_{t}$ to $c_{k}$ as -$s_{k}(z_{t})$. The resulting solution provides $p(z_{t}|c_{k})$, which we use in Eq. 7.
@@ 2.3 Indexing and Retrieval
We adopt the fast progressive search strategy proposed in our previous work [21]. Given a set $\mathcal{D}$ of audio tracks, each track $a_{i}\in\mathcal{D}$ is divided into overlapping segments of length $l$ s with a hop size of $h$ s. For each segment, we compute its representation $Z=\{z_{t}\}_{t=1}^{T}$, followed by its corresponding tokenized representation $Z^{\prime}=\{i_{t}\}_{t=1}^{T}$, where each $i_{t}$ denotes the index of the nearest codeword for $z_{t}$ in the codebook $C$, with $i_{t}\in\{1,2,\dots,K\}$. Subsequently, we construct a TF-IDF representation for each tokenized sequence $Z^{\prime}$ in the database and index them using IVF-PQ [9] for fast retrieval.
@@ 3.1 Databases
We trained the model on the LibriSpeech train-clean-360 subset and used the test-clean subset for validation. To construct the speech archive and queries, we selected the train-clean-100 subset, which contains approximately 100 hours of spoken content. This setup ensures that the evaluation is always conducted on speakers unseen during training our model. To assess cross-dataset generalization, we also evaluated our model on the TIMIT dataset [5]. We employed Montreal Forced Aligner [14] to obtain word alignments for both datasets. To emulate realistic acoustic conditions, the training data were augmented with a diverse range of background noise types and reverberation effects by incorporating noise samples and room impulse responses (RIRs) from the MUSAN corpus [22]. For evaluation, we employed RIRs from the Aachen Impulse Response Database [10] and noise clips from the ETSI background noise database [23], both of which contain recordings of real-world acoustic environments such as offices, public spaces, and street settings.
@@ 3.2 Evaluation
We evaluated the spoken term detection baselines using the Mean Term Weighted Value (MTWV) [19] as the primary performance metric. Also, to assess the robustness of speech tokenizers against speaker variations and audio distortions, we measured the Jaccard similarity [12] between the tokenized representations of different utterances of the same spoken term.
@@ 3.3 Baselines
We evaluated ASR-based baselines for STD tasks, including HuBERT [7] and WavLM [2], by extracting posterior tokens from these models. Also, we evaluated speech tokenizers that generate semantic speech tokens, such as WavLM, SpeechTokenizer [27], and BEST-STD [20], for STD tasks. For WavLM, we used the pretrained KMeans model available in the SpeechBrain toolkit [18], which was trained on features extracted from the 23rd layer of the large variant of WavLM. For SpeechTokenizer [27], we extracted the semantic tokens generated by the first quantizer in its residual vector quantization (RVQ) stack.
@@ 3.4 Implementation details
HuBERT-Large [7] 32 0.73 0.46 0.59 0.67 0.71 0.72 0.24 0.37 0.49 0.58 0.64
WavLM-Large [2] 32 0.72 0.62 0.67 0.70 0.71 0.71 0.52 0.60 0.65 0.68 0.70
@@
SpeechTokenizer [27] 1024 0.45 0.09 0.12 0.15 0.18 0.19 0.03 0.04 0.05 0.07 0.08
WavLM-Large [2] 1000 0.40 0.18 0.19 0.21 0.22 0.23 0.16 0.17 0.18 0.18 0.20
BEST-STD [20] 1024 0.72 0.21 0.29 0.42 0.60 0.65 0.19 0.22 0.38 0.55 0.62
@@ IV OOV IV OOV
HuBERT-Large [7] 0.13 0.21 0.30 0.40 0.47 0.47 0.16 0.27 0.34 0.40 0.41 0.43 0.14 0.22 0.31 0.43 0.49 0.51 0.16 0.28 0.37 0.43 0.44 0.46
WavLM-Large [2] 0.31 0.36 0.43 0.52 0.55 0.58 0.29 0.37 0.41 0.42 0.43 0.45 0.33 0.35 0.44 0.52 0.55 0.61 0.33 0.41 0.46 0.47 0.49 0.50
@@
SpeechTokenizer [27] 0.14 0.27 0.39 0.49 0.52 0.53 0.13 0.21 0.30 0.42 0.48 0.49 0.15 0.28 0.42 0.53 0.56 0.57 0.15 0.26 0.34 0.43 0.48 0.52
WavLM-Large [2] 0.17 0.34 0.40 0.53 0.55 0.55 0.17 0.25 0.35 0.43 0.47 0.49 0.19 0.38 0.44 0.57 0.59 0.61 0.19 0.29 0.35 0.46 0.47 0.51
BEST-STD [20] 0.27 0.35 0.43 0.50 0.57 0.62 0.22 0.29 0.37 0.44 0.49 0.54 0.29 0.38 0.47 0.54 0.62 0.66 0.25 0.33 0.40 0.49 0.50 0.56
@@
HuBERT-Large [7] 0.02 0.06 0.09 0.13 0.21 0.24 0.02 0.07 0.12 0.20 0.26 0.29 0.03 0.08 0.12 0.23 0.25 0.27 0.08 0.15 0.24 0.26 0.28 0.30
WavLM-Large [2] 0.11 0.18 0.24 0.30 0.32 0.36 0.15 0.22 0.29 0.31 0.35 0.37 0.12 0.21 0.23 0.35 0.37 0.39 0.18 0.24 0.31 0.35 0.39 0.41
@@
SpeechTokenizer [27] 0.03 0.05 0.11 0.14 0.18 0.20 0.02 0.04 0.06 0.11 0.13 0.16 0.05 0.12 0.18 0.19 0.23 0.23 0.07 0.11 0.14 0.18 0.21 0.23
WavLM-Large [2] 0.06 0.12 0.19 0.25 0.34 0.39 0.04 0.07 0.14 0.21 0.27 0.31 0.08 0.16 0.23 0.26 0.30 0.36 0.10 0.17 0.23 0.25 0.29 0.30
BEST-STD [20] 0.18 0.26 0.34 0.40 0.46 0.51 0.13 0.20 0.27 0.33 0.39 0.43 0.20 0.28 0.36 0.44 0.49 0.54 0.17 0.26 0.33 0.34 0.42 0.48
@@ 4.2 Codebook analysis
As shown in Table 3, our approach consistently achieves a normalized entropy close to 1, indicating near-perfect balance across different codebook sizes, ranging from 1024 to 4096. In contrast, KMeans-based tokenization [20] yields substantially lower entropy that drops further as the codebook size increases. Furthermore, we encountered codebook collapse when employing a trainable codebook regularized with KL divergence [1], despite extensive optimization efforts. These results demonstrate that our approach effectively overcomes the codebook collapse problem, enabling stable and scalable learning of large speech-token codebooks.
@@
KL Divergence [1] Learnable 0.63 0.50 0.38
BEST-STD [20] Kmeans 0.76 0.65 0.43
@@ 5 Conclusion
6 Acknowledgment
This research was supported by the Research Foundation Flanders (FWO) under grant S004923N of the SBO programme .